a boy jumps from a diving board at the pool. The board is 1.5 meters above the surface of the water. He bounces up in the air and then fall dowsn into the water. Exactly 2.0 seocnds pass from the time the boys' feet leave the baord to the time his feet hit the water
A) find the initial velocity
B) find the final velocity upon reaching the water
also answer to part a is +9.1m/s and answer to part b is -10.6m/s could you show me the approach

Answer by Alan3354(69443)

a boy jumps from a diving board at the pool. The board is 1.5 meters above the surface of the water. He bounces up in the air and then fall dowsn into the water. Exactly 2.0 seocnds pass from the time the boys' feet leave the baord to the time his feet hit the water
A) find the initial velocity
Using 9.8 m/sec/sec for gravity:
h(t) = -4.9t^2 + vt + 1.5 ---- v = speed leaving the board
h(2) = -4.9*2^2 + 2v + 1.5 = 0
-19.6 + 2v + 1.5 = 0
v = 9.05 m/sec
==================================
B) find the final velocity upon reaching the water
Speed is the 1st derivative of position.
h'(t) = -9.8t + v
h'(2) = -9.8*2 + 9.05
h'(2) = -10.55 m/sec